If Jose buys 15 apples and 20 oranges, his Php 500 is just enough. But if she buys 10 apples and 26 oranges, he will have Php 40 more for snacks. Find the cost of the apple.
Answer by mananth(16949)

Apples x unit price
Oranges y unit price

15 x + 20 y = 500 .............1
10 x + 26 y = 460 .............2
Eliminate y
multiply (1)by -26
Multiply (2) by 20
-390 x -520 y = -13000
200 x + 520 y = 9200
Add the two equations
-190 x = -3800
/ -190
x = 20
plug value of x in (1)
15 x + 20 y = 500
300 + 20 y = 500
20 y = 500 -300
20 y = 200
y = 10
Apples 20 unit price
Oranges 10 unit price
m.ananth@hotmail.ca